The simplified form of tan0/cot 0 is;
a. 1
b. (tan 0)^2
c. (cot 0)^2
d. (sin 0)(cos 0)^2

Answer by eperette(173)

Answer is b) (tan 0)^2
Two ways to look at it:
(1) Remember tan0 and cot0 are inverses of each other....in other words tan0=1/cot0 or cot0=1/tan0.....therefore tan0/cot0=tan0 / (1/tan0)= tan0 * tan0= (tan 0)^2....
(2) Or tan0=(sin0/cos0) and cot0=(cos0/sin0)....therefore tan0/cot0=(sin0/cos0) / (cos0/sin0)= (sin0/cos0)*(sin0/cos0)=(sin0)^2/(cos0)^2= (tan0)^2